OpenStudy (anonymous):

I need help with the set up in system of equation form How do I set this up? one solution containing 20 percent alcohol is mixed with another solution containing 60 percent alcohol to make 10 gallons of a solution that is 32 percent alcohol. How much of each solution is used?

OpenStudy (anonymous):

@Hoa

OpenStudy (anonymous):

x = amount of 20% alcohol y = amount of 60% alcohol eq. 1: x(20)+y(60)=(x+y)(32) but x+y = 10............. x = 10-y substitute this "y" in equation 1 x(2)+(10-y)(60)=(10)(32)

OpenStudy (anonymous):

follow?

OpenStudy (anonymous):

hold on let me see if i get it

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

kool check your answer.

OpenStudy (anonymous):

are the answers 7, and 3?

OpenStudy (anonymous):

and thanks hoa :) for helping

OpenStudy (anonymous):

I do nothing to help you, why thanks me?

OpenStudy (anonymous):

because you at least tryied by getting me helpo

OpenStudy (anonymous):

ok. I never say no when I can, if I cannot, i ask other to help. it's me. no need to sense